Product-Costing Accuracy, Plantwide and Departmental Rates, ABC

Escuha Company produces two type of calculators: scientific and business. Both products pass through two producing departments. The business calculator is by far the most popular. The following data have been gathered for these two products:

Product-Related Data
Scientific Business
Units produced per year 78,000 780,000
Prime costs $247,000 $2,470,000
Direct labor hours 96,500 965,000
Machine hours 53,000 530,000
Production runs 100 150
Inspection hours 2,000 3,000
Maintenance hours 2,250 9,000
Department Data
Department 1 Department 2
Direct labor hours:
  Scientific calculator 72,400 24,100
Business calculator 115,800 849,200
Total 188,200 873,300
Machine hours:
Scientific calculator 25,600 27,400
Business calculator 389,120 140,880
Total 414,720 168,280
Overhead costs:
Setup costs $214,000 $214,000
Inspection costs 165,000 165,000
Power 231,000 151,000
Maintenance 191,000 281,500
Total $801,000 $811,500

Required:

1. Compute the overhead cost per unit for each product using a plantwide, unit-based rate using direct labor hours. Round your calculations and answers to the nearest cent.

Plantwide rate $ per direct labor hour
Overhead cost per unit:
Scientific $ per unit
Business $ per unit

2. Compute the overhead cost per unit for each product using departmental rates. In calculating departmental rates, use machine hours for Department 1 and direct labor hours for Department 2. Round departmental rates and the overhead cost per unit to the nearest cent if rounding is required.

Departmental rates:
Department 1 $ per machine hour
Department 2 $ per direct labor hour
Overhead cost per unit:
Scientific $ per unit
Business $ per unit

Repeat using direct labor hours for Department 1 and machine hours for Department 2. Round departmental rates and the overhead cost per unit to the nearest cent if rounding is required.

Departmental rates:
Department 1 $ per direct labor hour
Department 2 $ per machine hour
Overhead cost per unit:
Scientific $ per unit
Business $ per unit

3. Compute the overhead cost per unit for each product using activity-based costing. Round your calculations and answers to the nearest cent.

Overhead Cost
Scientific $ per unit
Business $ per unit
Answer #1

Solution

Escuha Company

  1. Computation of the overhead cost per unit for each product using a plantwide, unit-based rate using direct labor hours:

Plantwide rate

$ per direct labor hour

Overhead cost per unit:

Scientific

$1.88 per unit

Business

$1.88 per unit

Computations:

Plantwide overhead rate = total overhead costs/direct labor hours

Total overhead cost = department 1 costs + department 2 costs

= $801,000 + 811,500 = $1,612,500

Total direct labor hours = department 1 DLH + department 2 DLH

= 188,200 + 873,300 = 1,061,500

Plandwide overhead rate = $1,612,500/1,061,500 = $1.52 per DLH

Overhead cost assigned = plantwide overhead rate x DLH used

Scientific Calculator –

Direct labor hours = 72,400 + 24,100 = 96,500

Overhead cost assigned = $1.52 x 96,500 = $146,591

Number of units = 78,000

Overhead cost per unit = $146,591/78,000 = $1.88

Business Calculator –

Direct labor hours = 965,000

Overhead cost assigned = $1.52 x 965,000 = $1,465,909

Number of units = 780,000

Overhead cost per unit = $1,465,909/780,000 = $1.88

  1. Computation of the overhead cost per unit for each product using a plantwide, unit based rate using department rates:

Department 1 – Machine hours; Department 2 – Direct labor hours

Department Rates

Department 1

$1.93 per MH

Department 2

$0.93 per DLH

Overhead Cost per Unit

Scientific

$0.92

Business

$1.97

Computations:

Department 1-

Total overhead costs = $801,000

Total machine hours = 414,720

Overhead rate per MH = $801,000/414,720 = $1.93

Department 1 overhead cost assigned to

Scientific Calculator –

MH = 25,600

Overhead rate = $1.93

Assigned overhead cost = $1.93 x 25,600 = $49,444

Number of units = 78,000

Department 1 overhead cost per unit = $0.63

Business Calculator –

MH = 389,120

Overhead rate = $1.93

Assigned overhead cost = $1.93 x 389,120 = $751,556

Number of units = 780,000

Department 1 overhead cost per unit = 751,556/780,000 = $0.96

Department 2-

Total overhead costs = $811,500

Total direct labor hours = 873,300

Overhead rate per DLH = $811,500/873,300 = $0.93

Department 2 overhead cost assigned to

Scientific Calculator –

DLH = 24,100

Overhead rate = $0.93

Assigned overhead cost = $0.93 x 24,100 = $22,395

Number of units = 78,000

Department 2 overhead cost per unit = 22,395/78,000 = $0.29

Business Calculator –

DLH = 849,200

Overhead rate = $0.93

Assigned overhead cost = $0.93 x 849,200 = $789,105

Number of units = 780,000

Department 2 overhead cost per unit = 789,105/780,000 = $1.012

Total overhead cost per unit –

Scientific Calculator -
department 1 = $0.63

Department 2 = $0.29

Total overhead cost per unit = $0.92

Business Calculator –

Department 1 = $0.96

Department 2 = $1.012

Total overhead cost per unit = $1.97

Computation of the overhead cost per unit for each product using a plantwide, unit-based rate using department rates:

Department 1 – Direct labor hours;          Department 2 – Machine hours

Department Rates

Department 1

$4.26 per DLH

Department 2

$4.82 per MH

Overhead Cost per Unit

Scientific

$5.64

Business

$1.50

Computations:

Department 1-

Total overhead costs = $801,000

Total direct labor hours = 188,200

Overhead rate per MH = $801,000/188,200 = $4.26

Department 1 overhead cost assigned to

Scientific Calculator –

DLH = 72,400

Overhead rate = $4.26

Assigned overhead cost = $4.26 x 72,400 = $308,142

Number of units = 78,000

Department 1 overhead cost per unit = 308,142/78,000 = $3.95

Business Calculator –

DLH = 115,800

Overhead rate = $4.26

Assigned overhead cost = $4.26 x 115,800 = $492,858

Number of units = 780,000

Department 1 overhead cost per unit = 492,858/780,000 = $0.63

Department 2-

Total overhead costs = $811,500

Total Machine hours = 168,280

Overhead rate per MH = $811,500/168,280 = $4.82

Department 2 overhead cost assigned to

Scientific Calculator –

MH = 27,400

Overhead rate = $4.82 per MH

Assigned overhead cost = $4.82 x 27,400 = $132,132

Number of units = 78,000

Department 2 overhead cost per unit = 132,132,78,000 = $1.69

Business Calculator –

MH = 140,880

Overhead rate = $4.82 per MH

Assigned overhead cost = $4.82 x 140,880 = $679,368

Number of units = 780,000

Department 2 overhead cost per unit = 679,368/780,000 = $0.87

Total overhead cost per unit –

Scientific Calculator -
department 1 = $3.95

Department 2 = $1.69

Total overhead cost per unit = $5.64

Business Calculator –

Department 1 = $0.63

Department 2 = $0.87

Total overhead cost per unit = $1.50

  1. Computation of the overhead cost per unit for each product using activity – based costing:

Overhead cost per unit

Scientific

$5.55

Business

$1.52

Computations:

Activity

Cost

Activity Usage

Activity rate = cost/usage

Setup

$428,000

250 production runs

$1,712 per production hour

Inspection

$330,000

5,000 inspection hours

$66 per inspection hour

Power

$382,000

583,000 machine hours

$0.66 per machine hour

Maintenance

$472,500

11,250 maintenance hours

$42 per maintenance hour

Scientific Calculator

Activity

Activity Rate

Activity Usage

Cost Assigned

Setup

$1,712 per production run

100 runs

$171,200

Inspection

$66 per inspection hour

2,000 inspection hours

$132,000

Power

$0.66 per MH

53,000 MH

$34,980

Maintenance

$42 per maintenance hour

2,250 maintenance hours

$94,500

Total

$432,680

number of units

78,000

Overhead cost per unit

$5.55

Business Calculator

Activity

Activity Rate

Activity Usage

Cost Assigned

Setup

$1,712 per production run

150 runs

$256,800

Inspection

$66 per inspection hour

3,000 inspection hours

$198,000

Power

$0.66 per MH

530,000 MH

$349,800

Maintenance

$42 per maintenance hour

9,000 maintenance hours

$378,000

Total

$1,182,600

number of units

780,000

Overhead cost per unit

$1.52